OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x - 5y if x = -5, y = 3 -5 5 35 -35

OpenStudy (anonymous):

4*-5 = -20 -5 * 3 = -15 -35 is the answerr

OpenStudy (anonymous):

simple. substituition

OpenStudy (anonymous):

4(-5) - 5(3) -20-15 -35